# AsynEIO: Asynchronous Monocular Event-Inertial Odometry Using Gaussian Process Regression

### v1

@@ VII Conclusion
This article introduces an asynchronous event-inertial odometry called AsynEIO that leverages a unified GP regression framework, where several important inertial fusion schemes are realized and evaluated. To the best of our knowledge, it is the first work that makes fair comparisons for asynchronous event-inertial fusion system. The expressing abilities of GP trajectories defined in different spaces and motion priors are assessed in detail, which can offer valuable suggest for choosing the suitable one according to certain conditions. The comparison analysis of various inertial factors explores their responses to sampling intervals and noises. We demonstrate that the GPP factor has greater contribution for inferring a fine-grained trajectory than GPIF and preintegration methods when the inertial measurements have low-level noises. Meanwhile, the GPP factor is more sensitive than other methods as the noise increases. Experiments conducted on real event-inertial sequences show that AsynEIO have a higher precision than the comparison method. Especially, our AsynEIO has a satisfied performance in high-speed maneuvering and low-illumination scenarios, which mainly benefited from the proposed asynchronous tracking and fusion mechanism. Although AsynEIO still need more improvements and engineering tricks for real-time estimation, we expect the proposed asynchronous event-inertial fusion framework and comparison results can contribute the community.
@@
VII-A Adjoint of $SE(3)$
$\bm{\mathcal{T}}=Ad(\exp(\bm{\xi}^{\land}))$ $=Ad(\begin{bmatrix}\bm{C}&\bm{C}\bm{J}\bm{\rho}\\ \bm{0}^{\top}&1\end{bmatrix})$
$=\begin{bmatrix}\bm{C}&\bm{0}\\ \bm{C}(\bm{J}\bm{\rho})^{\land}&\bm{C}\end{bmatrix}$
$=\exp(\bm{\xi}^{\curlywedge})$ (41)
where $\bm{C}=\exp(\bm{\phi}^{\land})$, $\bm{J}=\bm{J}(\bm{\phi})$ is the right Jacobian matrix.
$\bm{T}\exp(\bm{\xi}^{\land})\bm{T}^{-1}\equiv\exp((Ad(\bm{T})\bm{\xi})^{\land})$ (42)
VII-B Perturbing the Pose on $SE(3)$
$\bm{T}$ $=\bar{\bm{T}}\exp(\bm{\epsilon}^{\land})$ (43)
$\bm{T}$ $=\exp(\bm{\epsilon}^{\land})\bar{\bm{T}}$
$=\bar{\bm{T}}\exp((Ad(\bar{\bm{T}}^{-1})\bm{\epsilon})^{\land})$
$=\bar{\bm{T}}\exp((Ad(\bar{\bm{T}})^{-1}\bm{\epsilon})^{\land})$
$=\bar{\bm{T}}\exp(\bm{\mathcal{T}}^{-1}\bm{\epsilon})^{\land})$ (44)
VII-C Perturbing WNOJ Prior Residuals
@@
$=\log(\exp(-\bm{\epsilon}^{\land}_{k})\bar{\bm{T}}_{k}^{-1}\bar{\bm{T}}_{k+1}\exp(\bm{\epsilon}^{\land}_{k+1}))$
$\approx\bar{\bm{\xi}}_{k,k+1}+\bm{\mathcal{J}}^{-1}_{k,k+1,op}(\bm{\epsilon}_{k+1}-\exp(-\bm{\xi}^{\curlywedge})\bm{\epsilon}_{k})$
$=\bar{\bm{\xi}}_{k,k+1}+\underbrace{\bm{\mathcal{J}}_{k,k+1,op}^{-1}(\bm{\epsilon}_{k+1}-\bm{\mathcal{T}}_{k,k+1,op}^{-1}\bm{\epsilon}_{k})}_{\triangleq\delta\bm{\xi}_{k,k+1}}$ (45)
@@
$\approx\bm{\mathcal{J}}^{-1}(\bar{\bm{\xi}}_{k,k+1}+\bm{\mathcal{J}}_{k,k+1,op}^{-1}(\bm{\epsilon}_{k+1}-\bm{\mathcal{T}}_{k,k+1,op}^{-1}\bm{\epsilon}_{k}))$
$\approx\bm{\mathcal{J}}_{k,k+1,op}^{-1}+\frac{1}{2}(\bm{\mathcal{J}}_{k,k+1,op}^{-1}(\bm{\epsilon}_{k+1}-\bm{\mathcal{T}}_{k,k+1,op}^{-1}\bm{\epsilon}_{k}))^{\curlywedge}$
$=\bm{\mathcal{J}}_{k,k+1,op}^{-1}+\frac{1}{2}(\delta\bm{\xi}_{k,k+1})^{\curlywedge}$ (46)
$\bm{\varpi}_{k}$ $=\bar{\bm{\varpi}}_{k}+\delta\bm{\varpi}_{k}$ (47)
$\bm{\varpi}_{k+1}$ $=\bar{\bm{\varpi}}_{k+1}+\delta\bm{\varpi}_{k+1}$ (48)
$\dot{\bm{\varpi}}_{k}$ $=\bar{\dot{\bm{\varpi}}}_{k}+\delta\dot{\bm{\varpi}}_{k}$ (49)
$\dot{\bm{\varpi}}_{k+1}$ $=\bar{\dot{\bm{\varpi}}}_{k+1}+\delta\dot{\bm{\varpi}}_{k+1}$ (50)
@@
VII-D Jacobians of Interpolated State w.r.t Sampling States
@@
We first derive the perturbing Jacobians of $\bm{T}(\tau)$ w.r.t $\bm{T}_{k}$ and $\bm{\xi}_{k,\tau}$, respectively. Then, the perturbing Jacobians of $\bm{\xi}_{k,\tau}$ w.r.t $\bm{x}(k)$ and $\bm{x}(k+1)$ are given.
$\log(\bm{T}(\tau))^{\vee}$ $=\log(\bar{\bm{T}}(\tau)\exp(\bm{\epsilon}_{\tau}^{\land}))^{\vee}$
$\approx\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}\bm{\epsilon}_{\tau}$ (59)
@@
$=\log(\bm{T}_{k}\exp((\bar{\bm{\xi}}_{k,\tau}+\delta\bm{\xi}_{\tau})^{\land}))^{\vee}$
$=\log(\bm{T}_{k}\exp((\bar{\bm{\xi}}_{k,\tau})^{\land}))^{\vee}+\bm{\mathcal{J}}^{-1}\bm{\mathcal{J}}(\bar{\bm{\xi}}_{k,\tau})\delta\bm{\xi}_{\tau}$
$=\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}\bm{\mathcal{J}}(\bar{\bm{\xi}}_{k,\tau})\delta\bm{\xi}_{\tau}$ (60)
@@
$=\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}(\bm{\mathcal{T}}(\bm{\xi}_{k,\tau}))^{-1}\bm{\epsilon}_{k}$ (61)
@@
$=\bm{\mathcal{J}}(\bar{\bm{\xi}}_{k,\tau}+\delta\bm{\xi})(\bar{\dot{\bm{\xi}}}_{k,\tau}+\delta\dot{\bm{\xi}})$ (65)
$\approx\left(\bm{\mathcal{J}}(\bar{\bm{\xi}}_{k,\tau})-\frac{1}{2}\delta\bm{\xi}^{\curlywedge}\right)(\bar{\dot{\bm{\xi}}}_{k,\tau}+\delta\dot{\bm{\xi}})$ (66)
@@
$\frac{\partial\bm{\varpi}(\tau)}{\partial\bm{\xi}_{k,\tau}}$ $=\frac{1}{2}\bar{\dot{\bm{\xi}}}_{k,\tau}^{\curlywedge}$ (68)
$\frac{\partial\bm{\varpi}(\tau)}{\partial\dot{\bm{\xi}}_{k,\tau}}$ $=\bm{\mathcal{J}}_{k,\tau,op}$ (69)
$\frac{\partial\bm{\varpi}(\tau)}{\partial\ddot{\bm{\xi}}_{k,\tau}}$ $=\bm{0}$ (70)
@@
$\dot{\bm{\varpi}}(\tau)$ $=\bm{\mathcal{J}}_{k,\tau}\ddot{\bm{\xi}}_{k,\tau}+\frac{d}{dt}\bm{\mathcal{J}}_{k,\tau}\dot{\bm{\xi}}_{k,\tau}$
$\approx\bm{\mathcal{J}}_{k,\tau}\ddot{\bm{\xi}}_{k,\tau}-\frac{1}{2}\dot{\bm{\xi}}_{k,\tau}^{\curlywedge}\dot{\bm{\xi}}_{k,\tau}$
$=\bm{\mathcal{J}}_{k,\tau}\ddot{\bm{\xi}}_{k,\tau}.$ (71)
@@
$\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\bm{\xi}_{k,\tau}}$ $=\frac{1}{2}\bar{\ddot{\bm{\xi}}}_{k,\tau}^{\curlywedge}$ (72)
$\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\dot{\bm{\xi}}_{k,\tau}}$ $=\bm{0}$ (73)
$\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\ddot{\bm{\xi}}_{k,\tau}}$ $=\bm{\mathcal{J}}_{k,\tau,op}$ (74)
Perturbing the right-hand side of (56) with $\delta\bm{\varpi}_{k}$, $\delta\bm{\varpi}_{k+1}$ and $\delta\bm{\xi}_{k,k+1}$, we have
@@
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\bm{\epsilon}_{k}}$ $=-\frac{\partial\bm{\xi}_{k,\tau}}{\partial\bm{\epsilon}_{k+1}}\bm{\mathcal{T}}_{k,k+1,op}^{-1}$ (77)
@@
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\bm{\varpi}_{k}}$ $=\bm{\Lambda}_{12}$ (79)
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\dot{\bm{\varpi}}_{k}}$ $=\bm{\Lambda}_{13}$ (80)
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\dot{\bm{\varpi}}_{k+1}}$ $=\bm{\Psi}_{13}(\tau)\bm{\mathcal{J}}_{k,k+1,op}^{-1}$ (81)
VII-E Jacobians of Visual Residuals
$\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}^{\odot}$ $=\begin{bmatrix}-\bm{\kappa}^{\land}&\rho\bm{1}\\ \bm{0}^{\top}&\bm{0}^{\top}\end{bmatrix}$ (82)
$\bm{\xi}^{\land}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$ $\equiv\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}^{\odot}\bm{\xi}$ (83)
$\bm{T}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$ $=\left(\bm{T}_{1}^{-1}\bm{T}_{2}\right)^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$
$=\left((\bar{\bm{T}}_{1}\exp(\bm{\epsilon}_{1}^{\land}))^{-1}\bar{\bm{T}}_{2}\exp(\bm{\epsilon}_{2}^{\land})\right)^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$
$=\exp(-\bm{\epsilon}_{2}^{\land})\bar{\bm{T}}_{2}^{-1}\bar{\bm{T}}_{1}\exp(\bm{\epsilon}_{1}^{\land})\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$
$\approx\left(\bm{1}-\bm{\epsilon}_{2}^{\land}\right)\bar{\bm{T}}_{12}^{-1}\left(\bm{1}+\bm{\epsilon}_{1}^{\land}\right)\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$
$\approx\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}-\bm{\epsilon}_{2}^{\land}\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}+\bar{\bm{T}}_{12}^{-1}\bm{\epsilon}_{1}^{\land}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$
$=\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}-\left(\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}\right)^{\odot}\bm{\epsilon}_{2}+\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}^{\odot}\bm{\epsilon}_{1}$ (84)
$\bm{T}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$ $=\bm{T}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \bar{\rho}+\delta\rho\end{bmatrix}$
$=\bm{T}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \bar{\rho}\end{bmatrix}+\bm{T}_{12}^{-1}\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}\delta\rho$ (85)
$\bm{h}_{2}^{\dagger}=\bm{P}\bm{T}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}$ (86)
$\frac{\partial\bm{h}_{2}^{\dagger}}{\partial\bm{T}_{1}}$ $=\bm{P}\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}^{\odot}$ (87)
$\frac{\partial\bm{h}_{2}^{\dagger}}{\partial\bm{T}_{2}}$ $=-\bm{P}\left(\bar{\bm{T}}_{12}^{-1}\begin{bmatrix}\bm{\kappa}\\ \rho\end{bmatrix}\right)^{\odot}$ (88)
$\frac{\partial\bm{h}_{2}^{\dagger}}{\partial\rho}$ $=\bm{P}\bm{T}_{12}^{-1}\begin{bmatrix}\bm{0}\\ 1\end{bmatrix}$ (89)
$\frac{\partial\bm{q}_{2}}{\partial\bar{\bm{h}}_{2}}=\begin{bmatrix}f_{x}&0\\ 0&f_{y}\end{bmatrix}\begin{bmatrix}\frac{1}{h_{2,z}}&0&-\frac{h_{2,x}}{h_{2,z}^{2}}\\ 0&\frac{1}{h_{2,z}}&-\frac{h_{2,y}}{h_{2,z}^{2}}\\ \end{bmatrix}$ (90)
$\bm{T}_{1}$ $=\bm{T}_{ib1}\bm{T}_{bc}$ (91)
$\frac{\partial\bm{T}_{1}}{\partial\bm{T}_{ib1}}$ $=Ad(\bm{T}_{bc})^{-1}$ (92)

### v2

@@ VIII Conclusion
This article introduces an asynchronous event-inertial odometry called AsynEIO that leverages a unified GP regression framework, where several important inertial fusion schemes are realized and evaluated in the experiments. To the best of our knowledge, it is the first work that makes fair comparisons for asynchronous event-inertial fusion system. The expressing abilities of GP trajectories defined in different spaces and motion priors are assessed in detail, which can offer valuable suggestion for choosing the suitable one according to certain applications. The comparison analysis of various inertial factors demonstrates their properties. Experiments conducted on real event-inertial datasets with the state-of-the-art illustrate that AsynEIO has competitive accuracy and robustness. Especially, our AsynEIO has a satisfied performance in high-speed maneuvering and low-illumination scenarios, which mainly benefited from the proposed asynchronous tracking and fusion mechanism. Improving the time-efficiency of AsynEIO is an area of future work.
@@
VIII-A Perturbing the Pose on $SE(3)$
$\bm{T}$ $=\bar{\bm{T}}\exp(\bm{\epsilon_{1}}^{\land})=\exp(\bm{\epsilon_{2}}^{\land})\bar{\bm{T}}=\bar{\bm{T}}\exp((\bm{\mathcal{T}}^{-1}\bm{\epsilon_{2}})^{\land}),$ (40)
where $\bm{\mathcal{T}}=Ad(\exp(\bm{\xi}^{\land}))=\exp(\bm{\xi}^{\curlywedge})$ is the adjoint of $SE(3)$.
VIII-B Perturbing WNOJ Prior Residuals
@@
$\approx\bar{\bm{\xi}}_{k,k+1}+\underbrace{\bm{\mathcal{J}}_{k,k+1,op}^{-1}(\bm{\epsilon}_{k+1}-\bm{\mathcal{T}}_{k,k+1,op}^{-1}\bm{\epsilon}_{k})}_{\triangleq\delta\bm{\xi}_{k,k+1}}$ (41)
@@
$\approx\bm{\mathcal{J}}_{k,k+1,op}^{-1}+\frac{1}{2}(\delta\bm{\xi}_{k,k+1})^{\curlywedge}$ (42)
@@
VIII-C Jacobians of Interpolated State w.r.t Sampling States
@@
We first derive the perturbing Jacobians of $\bm{T}(\tau)$ w.r.t $\bm{T}_{k}$ and $\bm{\xi}_{k,\tau}$, respectively. Then, the perturbing Jacobians of $\bm{\xi}_{k,\tau}$ w.r.t $\bm{x}(k)$ and $\bm{x}(k+1)$ are given by
$\log(\bm{T}(\tau))^{\vee}$ $=\log(\bar{\bm{T}}(\tau)\exp(\bm{\epsilon}_{\tau}^{\land}))^{\vee}\approx\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}\bm{\epsilon}_{\tau}$ (51)
@@
$\approx\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}\bm{\mathcal{J}}(\bar{\bm{\xi}}_{k,\tau})\delta\bm{\xi}_{\tau}$ (52)
@@
$\approx\log(\bar{\bm{T}}(\tau))^{\vee}+\bm{\mathcal{J}}^{-1}(\bm{\mathcal{T}}(\bm{\xi}_{k,\tau}))^{-1}\bm{\epsilon}_{k}$ (53)
@@
$\frac{\partial\bm{\varpi}(\tau)}{\partial\bm{\xi}_{k,\tau}}=\frac{1}{2}\bar{\dot{\bm{\xi}}}_{k,\tau}^{\curlywedge},\quad\frac{\partial\bm{\varpi}(\tau)}{\partial\dot{\bm{\xi}}_{k,\tau}}=\bm{\mathcal{J}}_{k,\tau,op},\quad\frac{\partial\bm{\varpi}(\tau)}{\partial\ddot{\bm{\xi}}_{k,\tau}}$ $=\bm{0}$ (58)
@@
$\dot{\bm{\varpi}}(\tau)$ $=\bm{\mathcal{J}}_{k,\tau}\ddot{\bm{\xi}}_{k,\tau}+\frac{d}{dt}\bm{\mathcal{J}}_{k,\tau}\dot{\bm{\xi}}_{k,\tau}$ (59)
@@
$\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\bm{\xi}_{k,\tau}}=\frac{1}{2}\bar{\ddot{\bm{\xi}}}_{k,\tau}^{\curlywedge},\quad\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\dot{\bm{\xi}}_{k,\tau}}=\bm{0},\quad\frac{\partial\dot{\bm{\varpi}}(\tau)}{\partial\ddot{\bm{\xi}}_{k,\tau}}=\bm{\mathcal{J}}_{k,\tau,op}$ (60)
Perturbing the right-hand side of (VIII-C) with $\delta\bm{\varpi}_{k}$, $\delta\bm{\varpi}_{k+1}$ and $\delta\bm{\xi}_{k,k+1}$, we have
@@
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\bm{\epsilon}_{k}}$ $=-\frac{\partial\bm{\xi}_{k,\tau}}{\partial\bm{\epsilon}_{k+1}}\bm{\mathcal{T}}_{k,k+1,op}^{-1},\quad\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\bm{\varpi}_{k}}=\bm{\Lambda}_{12}$ (63)
@@
$\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\dot{\bm{\varpi}}_{k}}$ $=\bm{\Lambda}_{13}\quad\frac{\partial\bm{\xi}_{k,\tau}}{\partial\delta\dot{\bm{\varpi}}_{k+1}}=\bm{\Psi}_{13}(\tau)\bm{\mathcal{J}}_{k,k+1,op}^{-1}$ (65)